OpenStudy (anonymous):

lim x-> -infinity x^2e^x

OpenStudy (zzr0ck3r):

2*e^(-infinity) = 2*e^(1/infinity) = 2e^0 = 2 -infinity^2 = infinity

OpenStudy (zzr0ck3r):

get it @juliweaver ?

OpenStudy (anonymous):

not really sure

OpenStudy (zzr0ck3r):

do you understand 1/-infinity = 0?

OpenStudy (zzr0ck3r):

do you understand e^(-x) = e^(1/x)

OpenStudy (anonymous):

how are you getting to 1/-infinity? Is it because e raised to any power is 0?

OpenStudy (anonymous):

I get your last statement yes

OpenStudy (zzr0ck3r):

ok if x^(2e^(-x)) = x^(2e^(1/x)) so if we plug in -infinity we get (-infinity)^(2e^(1/-infinity) 1/-infinity = 0 so we have (-infinity)^(2*e^0) = (-infinity)^(2*1) = (-infinity)^2

OpenStudy (anonymous):

okay their should be a space x^2 e^x

OpenStudy (anonymous):

to the negative infinity

OpenStudy (zzr0ck3r):

\[x^{2} * e^{x} ?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so using l'hopitals rule how do you solve?

OpenStudy (anonymous):

why?

OpenStudy (zzr0ck3r):

one sec

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

would it be -infinity?